Why does it requires great amount of heat to increase the temperature of 10kg water from 10° to 20°


1
Expert's answer
2021-05-26T10:31:17-0400

Because water has great specific heat capacity "c= 4200\frac {J}{kg\cdot °C}" and

"Q=cm \Delta t=4200\cdot 10\cdot (20-10)=0.42~MJ."